a) Explain the equation 𝑭 = −𝜵𝑈


1
Expert's answer
2021-02-12T16:17:34-0500

"dA=\vec{F}d\vec{r}=-dU,"

"\vec{F}=F_x\vec{i}+F_y\vec{j}+F_z\vec{k},"

"d\vec{r}=dx\vec{i}+dy\vec{j}+dz\vec{k},"

"\vec{F}d\vec{r}=F_xdx+F_ydy+F_zdz=-dU,"

"F_x\partial x=-\partial U,~" "F_x=-\frac{\partial U}{\partial x},"

"F_y=\frac{\partial U}{\partial y}," "F_z=\frac{\partial U}{\partial z},"

"\vec{F}=-(\frac{\partial U}{\partial x}\vec{i}+\frac{\partial U}{\partial y}\vec{j}+\frac{\partial U}{\partial z}\vec{k}),"

"\vec{F}=-\text{grad}U,"

"\vec{F}=-\nabla U."